1) Life insurance salesman meets his prospective customers separately. At each meeting, he makes the same effort to persuade the prospective client to be insured. The vendor's understanding after a long term in the insurance business is that the probability of persuading the customer to be insured (probability of success) is 0,1. What is the probability in 4 trials to insure one of them?
1
Expert's answer
2020-01-15T08:55:39-0500

As probability of success p = 0.1,


Probability of failure q = 1 - p


= 1 - 0.1 = 0.9

Number of trials = 4

Number of success = 1

Now formula for binomial distribution with x success out of n trials.

P(x) = nCx * px * qn-x

P(x = 1) = 4C1 * 0.11 * 0.93 = 0.2916